OpenStudy (anonymous):

what is the square root of 1.01 using 5 as exponent

OpenStudy (dumbcow):

i dont understand the question sqrt(1.01) = 1.00499

OpenStudy (anonymous):

i need to evaluate the problem 5/1.01. (i dont have a squaare root sign so i will use this. I need to know what 5 numbers will add up to 1.01, which will be the squre root of that number

OpenStudy (anonymous):

\[\sqrt[5]{1.01?}\]

OpenStudy (anonymous):

evaluate \[\sqrt[5]{1.01}\]

OpenStudy (anonymous):

Not sure what the problem is either. The equation below says five times a number is equal to the square root of 1.01: \[5 x=\sqrt{\frac{101}{100}} \] The answer for x is \[\frac{\sqrt{101}}{50} \text{or } 0.200998 \] \[5*0.200998 =1.00499 \]